OpenStudy (anonymous):

Coach Johnson is 10 years older than Tim. When Tim started football 3 years ago, Coach Johnson was twice his age. How old is Tim now?

OpenStudy (anonymous):

let the age of tim and caoch be x and x+10 by the problem (x+10-3)=2(x-3) or x+7=2x-6 or x=6+7 or x=13 hence tim,s age now is 13